# Beyond Ensembles: Simulating All-Atom Protein Dynamics in a Learned Latent Space

### v3

@@ Abstract
Simulating the long-timescale dynamics of biomolecules is a central challenge in computational science. While enhanced sampling methods can accelerate these simulations, they rely on pre-defined collective variables that are often difficult to identify. A recent generative model, LD-FPG, demonstrated that this problem could be bypassed by learning to sample the static equilibrium ensemble as all-atom deformations from a reference structure, establishing a powerful method for all-atom ensemble generation. However, while this approach successfully captures a system’s probable conformations, it does not model the temporal evolution between them. We introduce the Graph Latent Dynamics Propagator (GLDP), a modular component for simulating dynamics within the learned latent space of LD-FPG. We then compare three classes of propagators: (i) score-guided Langevin dynamics, (ii) Koopman-based linear operators, and (iii) autoregressive neural networks. Within a unified encoder–propagator–decoder framework, we evaluate long-horizon stability, backbone and side-chain ensemble fidelity, and functional free-energy landscapes. Autoregressive neural networks deliver the most robust long rollouts; score-guided Langevin best recovers side-chain thermodynamics when the score is well learned; and Koopman provides an interpretable, lightweight baseline that tends to damp fluctuations. These results clarify the trade-offs among propagators and offer practical guidance for latent-space simulators of all-atom protein dynamics.
1 Signal Processing Laboratory (LTS2), EPFL, Lausanne, Switzerland
2 Institute of Bioengineering, EPFL, Lausanne, Switzerland
3 Ludwig Institute for Cancer Research, Lausanne, Switzerland
Corresponding author: aditya.sengar@epfl.ch
@@ 1 Introduction
Molecular simulations are indispensable for studying the complex dynamics that govern biological function, yet brute-force approaches struggle to access the slow, functionally relevant motions—such as protein folding, ligand binding, or allosteric switching—due to rugged energy landscapes and the dominance of rare events [1, 2]. To mitigate this gap—beyond what enhanced sampling can offer when suitable collective variables are hard to specify  [3] —a complementary strategy has gained traction: recasting the simulation challenge from one of brute-force integration to one of representation learning. In this representation-first view, the simulation problem is recast as a modular encoder–propagator–decoder pipeline: an encoder maps high-dimensional atomic configurations into a continuous, low-dimensional latent space; a propagator evolves the system’s state within this simplified space; and a decoder maps the resulting latent trajectory back to all-atom coordinates [4, 5].
Progress in this paradigm follows two complementary directions: The first focuses on learning the underlying physics, employing score-based diffusion  [6, 7], flow matching  [8], and energy-based models  [9] to learn generative surrogates—sometimes yielding differentiable force fields—that implicitly define the system’s potential of mean force [10, 11, 12, 13]. The second centers on learning simplified dynamical coordinates, using time-aware autoencoders or Koopman/DMD analysis  [14, 15] to discover an intrinsic manifold where the long-term dynamics become stable, predictable, or even approximately linear [16, 17, 18, 19, 20].
Despite the promise of latent-space simulation, the choice of propagator—the engine that advances the state—entails a trade-off between physical faithfulness, long-horizon stability, and expressive capacity. We adopt LD-FPG [21] as a fixed encoder–decoder and compare, in the same latent, three propagators: (i) score-guided Langevin dynamics, (ii) a Koopman linear operator, and (iii) an autoregressive neural network (NN). Our analysis reveals a clear trade-off: the MLP is most stable, the score-guided model is most physically detailed for local motions, and the Koopman operator is a simple but overly rigid baseline.
@@
(1) A controlled study that holds the LD-FPG encoder–decoder fixed while swapping the latent propagator, isolating where gains arise.(2) Three propagators within one latent: score-guided Langevin, a Koopman operator, and an autoregressive NN.(3) Benchmarks on alanine dipeptide, 7JFL, and $_{\mbox{1}}A$R.(4) First learned, all-atom rollout (to our knowledge) that recovers a GPCR activation surface ($_{\mbox{2}}A$R).
@@ 2 Related Work
The encoder–propagator–decoder blueprint.
@@
Simulating in a low-dimensional space is well established. Molecular Latent Space Simulators (LSS) [4] and Deep Generative MSMs (DeepGenMSM) [5] factorize the task into an encoder that maps coordinates to latent variables, a propagator that advances them, and a decoder that returns all-atom structures. Realizations include autoregressive simulators with recurrent propagators [22], trajectory-level generators that synthesize MD clips in coordinate space [23], LSS variants that target free-energy surfaces and transition kinetics [24], and joint conformation–dynamics models trained via autoregression [25]. We follow this modular design but (i) hold the LD-FPG decoder fixed to control reconstruction quality and (ii) focus our analysis on the choice of latent propagator.
Learning dynamically aware latent spaces (encoders).
The quality of a latent simulation hinges on the quality of the latent space itself. While linear methods like time-lagged independent component analysis (TICA) remain strong baselines for identifying slow variables [26], deep learning has enabled far more expressive encoders. Time-lagged autoencoders such as the Variational Dynamics Encoder (VDE) learn non-linear representations predictive over a time delay $\Delta t$ [16, 27]. VAMPnets use a variational principle to approximate the leading Koopman eigenfunctions, corresponding to the slowest dynamical processes [17]. To remove hand-crafted features, modern approaches leverage graph neural networks (GNNs) to learn CVs directly from coordinates in a permutation- and symmetry-aware manner [28, 20, 29], often with information-bottleneck objectives to explicitly optimize predictiveness of future states [30]. For a recent overview of time-lagged encoders and related autoencoder variants, see [31].
Linear dynamics in latent space (Koopman/DMD propagators).
The Koopman-operator framework provides a route to linearize non-linear dynamics in a learned observable space [18]. Extended dynamic mode decomposition (EDMD) [32] and DMD [15] showed that a wide class of systems admits accurate linear predictors; subsequent work learns such observables end-to-end with deep encoders [14]. This has inspired Koopman autoencoders that enforce a linear evolution rule within the latent space [19, 33, 34], often yielding strong long-horizon stability by reducing compounding errors in nonlinear rollouts. Beyond linear Koopman rules, implicit transfer-operator learning uses a diffusion model to parameterize the Markov kernel at multiple lags, yielding consistent long-horizon rollouts under SE(3) symmetry  [35].
@@
Non-linear sequence models (neural propagators).
A complementary path directly learns the non-linear transition function. The Learning Effective Dynamics (LED) framework uses LSTMs to propagate latent variables and capture memory effects [22]; others pair RNNs with physical resampling such as Maximum Caliber to enforce kinetic consistency [36]. Continuous-time learners (Neural ODEs) offer flexible parametric flows [37], with extensions to stochastic settings via Neural SDEs [38]. Recent work also couples an autoregressive temporal module with an SE(3)-equivariant diffusion decoder to jointly learn conformations and their ordering from MD (ConfRover)  [25]. Our MLP propagator serves as a streamlined baseline in this family.
Physics-guided stochastic models (Langevin and diffusion).
Generative models increasingly embed statistical-mechanics structure. Score-based diffusion has been used to learn effective force fields for coarse-grained MD [10], tying the denoising objective to forces and Langevin-type simulators via the SDE view of score models [6]. For accelerated evolution, frame-to-frame diffusion/score surrogates have been proposed, including DiffMD for trajectory synthesis conditioned on the previous frame [39] and Score Dynamics, which targets picosecond time steps while preserving short-time kinetic signals [12]. Flow-matching has likewise been applied to coarse-grained rollouts without explicit forces (F3low; manifold/SE(3)-equivariant) [13] and to learning coarse-grained free-energy models via density estimation [40]. A recurring theme is consistency between the sampled equilibrium ensemble and the stationary law of the learned dynamics; enforcing Fokker–Planck consistency during training improves long-horizon behavior [11].
GPCRs as a proving ground for slow dynamics.
G protein-coupled receptors (GPCRs) are a demanding testbed for dynamic models, as their function is governed by slow transitions between metastable states [1, 2]. Decades of structural and computational work have characterized the micro-switches and allosteric pathways involved in their activation [41, 42], making them ideal systems for benchmarking. Our evaluation on $_{\mbox{1}}A$R and $_{\mbox{2}}A$R tests the ability of each latent propagator to recover these functionally critical motions.
Positioning among recent generative simulators.
While geometric diffusion models achieve strong results for static 3D structures and poses [43, 44], our focus is the distinct problem of simulating temporal dynamics. Recent efforts include coordinate-space trajectory synthesis with diffusion (GeoTDM) [45], SE(3)-equivariant flow matching for frame-to-frame coarse-grained rollouts (F3low) [13], and neural operators for multi-step 3D dynamics (EGNO) [46]. In parallel, latent and physics-informed approaches aim to accelerate sampling or encode structure—LAST for adaptive MD [47], ConfRover for joint conformation–dynamics learning [25], and NeuralMD, a multi-grained group-symmetric differential-equation model for protein–ligand binding dynamics [48]. Complementing these directions, we isolate the propagator by comparing linear (Koopman), neural, and score-guided dynamics within the same learned latent and fixed decoder, enabling a controlled comparison of long-horizon stability, ensemble fidelity, and functional free-energy surfaces.
@@ 3.1 GLDP: Propagation in the LD-FPG Latent Space
We adopt the LD-FPG encoder–decoder backbone. The encoder, a Chebyshev Graph Neural Network (ChebNet), maps all-atom coordinates $X(t)\in\mathbb{R}^{N\times 3}$ to a pooled latent $z(t)\in\mathbb{R}^{d}$. Trained on an MD trajectory, the encoder produces a time series $\{z_{0},z_{1},\dots,z_{M}\}$ that captures conformational variation. We then learn a latent-time update of the form
@@ 3.4 Score-Guided Langevin Propagator
Our propagators evolve only the latent coordinate $z$ (no momenta, no barostat/box dynamics). The objective is to reproduce the configuration-space equilibrium of the reference MD at its thermodynamic state. Let $p_{X}(x)$ denote the equilibrium distribution of all-atom coordinates $x$ from the source MD, and let $E[\cdot]$ and $D[\cdot]$ be the fixed LD-FPG encoder and decoder. The target latent law is the pushforward $p_{Z}\;=\;E_{\#}\,p_{X}\quad\text{with}\quad z=E(x)$ and decoded samples are $\hat{x}=D(z)$. We fit propagators so that the stationary decoded trajectory $\{\hat{x}(t)\}$ matches reference statistics in $X$-space (e.g., dihedral histograms, RMSF, contacts, and TM-distance free-energy surfaces). Because no Hamiltonian with momenta is integrated in latent space, we do not claim NVE energy conservation or exact physical kinetics. The score-guided Langevin propagator targets a canonical law in latent space whose invariant density is $p_{Z}$ when the score is exact; Koopman and the autoregressive NN model approximate the one-step transition kernel $K(z_{t+1}\!\mid z_{t})$ observed from $(z_{t},z_{t+1})$ under the fixed encoder. Unless stated otherwise, comparisons of timescales are diagnostic rather than calibrated to wall-clock time.
@@ 4 Results and Discussion
Key result. GLDP operates reliably across scales (ADP $\rightarrow$ 7JFL $\rightarrow$ $_{\mbox{1}}A$R) and reproduces the $_{\mbox{2}}A$Ractivation surface; this cross-scale pattern is the main message of the paper. See Appendix A.7 for $_{\mbox{1}}A$R ablations (encoder width, pooling size, and Langevin/denoiser settings).
@@ 4.1 Comparative Benchmarks on Trajectory Generation
To benchmark GLDP, we compared it against a diverse set of state-of-the-art models: the conceptually similar Latent Space Simulator (LSS) [4], the video-synthesis-based MD-Gen [23], and GeoTDM [45], a leading coordinate-based diffusion model. All methods were evaluated on equilibrium fidelity, flexibility, and kinetics across systems of increasing complexity: the canonical alanine dipeptide (ADP), the 7JFL protein, and the functionally complex $_{\mbox{1}}A$R GPCR. The quantitative results are summarized in Table 1.  [^1] GeoTDM on 7JFL and $_{\mbox{1}}A$R exceeded 80 GB GPU memory (H100). MD-Gen does not support capped termini in our ADP setup; ADP results omitted.
Table 1: Benchmark of dynamics models. GLDP is compared against LSS, MD-Gen, and GeoTDM. Lower JSD is better. $B$: backbone; SC: sidechain. For RMSF, predicted values closer to the ground-truth row are better. Higher contact correlation is better. Bold indicates the best result among models. N/A: not applicable.
JSD Avg. RMSF (Å)
System Model Dihedral (BB/SC) Coord. (BB/SC) Contact Corr. Backbone Sidechain
ADP Ground Truth N/A N/A N/A 0.56 1.34
LSS 0.016 / N/A 0.621 / 0.633 N/A 0.54 0.90
GeoTDM 0.123 / N/A 0.444 / 0.395 N/A 0.43 0.77
GLDP 0.061 / N/A 0.366 / 0.386 N/A 0.61 0.89
7JFL Ground Truth N/A N/A N/A 2.09 3.05
LSS 0.027 / 0.097 0.691 / 0.682 0.935 0.60 0.78
MD-Gen 0.035 / 0.090 0.377 / 0.361 0.960 0.99 1.73
GLDP 0.037 / 0.081 0.426 / 0.459 0.949 1.44 2.30
$_{\mbox{1}}A$R Ground Truth N/A N/A N/A 1.49 1.93
LSS 0.134 / 0.146 0.657 / 0.602 0.971 1.11 1.41
MD-Gen 0.033 / 0.088 0.106 / 0.117 0.941 0.86 1.31
GLDP 0.019 / 0.067 0.063 / 0.087 0.985 1.22 1.48
@@
We assess equilibrium fidelity via Jensen–Shannon divergence (JSD) over global coordinates and local dihedrals. On the small ADP system, GLDP attains the best coordinate JSD, while LSS edges out on backbone dihedral JSD and is closest to ground-truth RMSF. On the medium 7JFL system, MD-Gen leads on coordinate JSD and contact correlation, LSS is best on backbone-dihedral JSD, and GLDP is best on side-chain–dihedral JSD and most closely matches backbone/side-chain RMSF amplitudes. On the large GPCR $_{\mbox{1}}A$R, GLDP achieves the lowest JSDs for both coordinates and dihedrals, the highest contact correlation, and the closest RMSFs, indicating the most faithful recovery of both tertiary structure and local conformational statistics. Overall, GLDP provides the strongest balance across metrics on complex proteins, while MD-Gen and LSS retain advantages on specific axes for smaller or mid-scale systems.
Molecular flexibility (RMSF).
Matching the amplitude of motion is as important as matching mean structure, so we compare backbone/side-chain RMSF to the reference MD. On ADP, all methods undershoot side-chain flexibility; GLDP slightly overestimates backbone RMSF (0.61 Å vs. 0.56 Å), while LSS is the closest overall on ADP (0.54/0.90 Å). On 7JFL, GLDP best recovers fluctuation magnitudes for both backbone and side chains (1.44/2.30 Å vs. 2.09/3.05 Å), whereas MD-Gen (0.99/1.73 Å) and LSS (0.60/0.78 Å) are overly rigid. On $_{\mbox{1}}A$R, all models are conservative, but GLDP remains closest to the reference variability (1.22/1.48 Å vs. 1.49/1.93 Å). In sum, GLDP most faithfully reproduces fluctuation amplitudes on the larger proteins, complementing its strong ensemble statistics.
@@
As a coarse proxy for long-timescale stability, we compare the Pearson correlation between model and reference time-averaged contact maps (Appendix A.6). Contact patterns are well preserved by the top models (Pearson $r>0.93$ where evaluated). On 7JFL, MD-Gen best matches the contacts ($r=0.960$), while GLDP attains the highest correlation on $_{\mbox{1}}A$R ($r=0.985$) (ADP omitted: system too small for a meaningful contact map.)
@@
Across flexibility and kinetics, GLDP offers an excellent balance of realism and stability. On the complex $_{\mbox{1}}A$R receptor, it is the top performer across all metrics, achieving the lowest Jensen-Shannon Divergence for both coordinates and dihedrals, the highest contact correlation ($r=0.985$), and the most accurate fluctuation amplitudes (RMSF). This strong performance extends to other systems, where GLDP consistently places as either the best-performing model—such as for coordinate fidelity on ADP and side-chain distributions on 7JFL—or as a close competitor to the leader. Crucially, GLDP faithfully matches the flexibility of larger proteins where competing methods are often overly rigid, positioning it as a highly dependable latent-space simulator.

### v4

@@ Abstract
Simulating the long-timescale dynamics of biomolecules is a central challenge in computational science. While enhanced sampling methods can accelerate these simulations, they rely on pre-defined collective variables that are often difficult to identify, restricting their ability to model complex switching mechanisms between metastable states. A recent generative model, LD-FPG, demonstrated that this problem could be bypassed by learning to sample the static equilibrium ensemble as all-atom deformations from a reference structure, establishing a powerful method for all-atom ensemble generation. However, while this approach successfully captures a system’s probable conformations, it does not model the temporal evolution between them. We introduce the Graph Latent Dynamics Propagator (GLDP), a modular component for simulating dynamics within the learned latent space of LD-FPG. We then compare three classes of propagators: (i) score-guided Langevin dynamics, (ii) Koopman-based linear operators, and (iii) autoregressive neural networks. Within a unified encoder–propagator–decoder framework, we evaluate long-horizon stability, backbone and side-chain ensemble fidelity, and temporal kinetics via TICA. Benchmarks on systems ranging from small peptides to mixed-topology proteins and large GPCRs reveal that autoregressive neural networks deliver the most robust long rollouts and coherent physical timescales; score-guided Langevin best recovers side-chain thermodynamics when the score is well learned; and Koopman provides an interpretable, lightweight baseline that tends to damp fluctuations. These results clarify the trade-offs among propagators and offer practical guidance for latent-space simulators of all-atom protein dynamics.
@@ 1 Introduction
Molecular simulations are indispensable for studying the complex dynamics that govern biological function, yet brute-force approaches struggle to access the slow, functionally relevant motions—such as protein folding, ligand binding, or allosteric switching—due to rugged energy landscapes and the dominance of rare events [9, 21]. To mitigate this gap, beyond what enhanced sampling can offer when suitable collective variables are hard to specify  [5], a complementary strategy has gained traction: recasting the simulation challenge from one of brute-force integration to one of representation learning. In this representation-first view, the simulation problem is recast as a modular encoder–propagator–decoder pipeline: an encoder maps high-dimensional atomic configurations into a continuous, low-dimensional latent space; a propagator evolves the system’s state within this simplified space; and a decoder maps the resulting latent trajectory back to all-atom coordinates [35, 48].
Progress in this paradigm follows two complementary directions: The first focuses on learning the underlying physics, employing score-based diffusion  [36, 14], flow matching  [24], and energy-based models  [22] to learn generative surrogates (sometimes yielding differentiable force fields) that implicitly define the system’s potential of mean force [1, 30, 15, 23]. The second centers on learning simplified dynamical coordinates, using time-aware autoencoders or Koopman/DMD analysis  [26, 40] to discover an intrinsic manifold where the long-term dynamics become stable, predictable, or even approximately linear [13, 27, 3, 2, 11].
Despite the promise of latent-space simulation, the choice of propagator (the engine that advances the state) entails a trade-off between physical faithfulness, long-horizon stability, and expressive capacity. We adopt LD-FPG [33] as a fixed encoder–decoder and compare, in the same latent, three propagators: (i) score-guided Langevin dynamics, (ii) a Koopman linear operator, and (iii) an autoregressive neural network (NN). Our analysis reveals a clear trade-off: Autoregressive NN is most stable, the score-guided model is most physically detailed for local motions, and the Koopman operator is a simple but overly rigid baseline.
GLDP is designed as a system-specific surrogate for MD, akin to LSS, GeoTDM, ITO [35, 12, 32]. The objective is to learn the specific transition operator of a target system to enable accelerated sampling of its equilibrium and kinetics. While cross-protein generalization (as pursued by foundation models) is a valuable complementary goal, this work focuses on maximizing thermodynamic fidelity and long-horizon stability for specific, complex targets.
@@
(1) A controlled study that holds the LD-FPG encoder–decoder fixed while swapping the latent propagator, isolating where gains arise. (2) Three propagators within one latent: score-guided Langevin, a Koopman operator, and an autoregressive NN. (3) Benchmarks on alanine dipeptide, 7JFL, HIV-1 Protease (1R6W), and $_{\mbox{1}}A$R. (4) First learned, all-atom rollout (to our knowledge) that recovers a GPCR activation surface ($_{\mbox{2}}A$R).
@@ 2 Related Work
Latent simulators and encoder–propagator–decoder pipelines.
@@
Simulating dynamics in a low-dimensional latent space has been explored in several molecular simulators. Molecular Latent Space Simulators (LSS) [35] and Deep Generative MSMs (DeepGenMSM) [48] already factorize the task into an encoder that maps coordinates to latent variables, a propagator that advances them, and a decoder that returns all-atom structures. Follow-up work uses recurrent propagators [42], trajectory-level generators that synthesize MD clips in coordinate space [16], and LSS variants that target free-energy surfaces and transition kinetics [8]. Joint conformation–dynamics models trained via autoregression have also been proposed [34]. Our framework follows this modular design but differs in two respects: (i) we freeze the LD-FPG encoder–decoder [33] to control reconstruction quality and (ii) we perform a focused comparison of different propagators within the same latent, isolating how the propagation rule affects stability, ensemble fidelity, and functional observables.
Learning dynamically aware latent spaces.
The quality of a latent simulator is largely determined by its encoder. Classical TICA is a standard tool for discovering slow coordinates [29], while deep time-lagged autoencoders such as VDE learn nonlinear, delay-predictive representations [13, 45]. VAMPnets approximate leading Koopman eigenfunctions to capture the slowest processes [27], and more recent work uses GNNs to learn symmetry-aware collective variables directly from coordinates, sometimes with information-bottleneck objectives to optimize predictiveness of future states [52, 11, 51, 44]; see Chen & Chipot [4] for a survey. In this work we instead keep the LD-FPG ChebNet encoder fixed [33] and treat its latent as given, focusing on how different propagation mechanisms behave once a reconstruction-ready latent space has been trained.
@@
Propagators: linear Koopman, neural sequence models, and stochastic dynamics.
The Koopman-operator view approximates nonlinear dynamics with a linear map in a learned observable space [3], with EDMD/DMD estimating this map directly [47, 40] and deep variants learning the observables so that latent evolution is linear [26, 2, 37, 28]; implicit transfer-operator learning further parameterizes Markov kernels with diffusion models under SE(3) symmetry [32]. A second line directly fits nonlinear transition maps using neural sequence models, including LSTMs/RNNs for effective dynamics [42], RNNs with Maximum Caliber constraints [39], and continuous-time Neural ODE/SDE formulations [6, 18], with ConfRover combining an autoregressive temporal module and an SE(3)-equivariant diffusion decoder [34]. A third class uses stochastic dynamics guided by generative models: score-based diffusion linked to Langevin-type simulators and effective force fields [36, 1, 15], and flow matching for coarse-grained rollouts and free-energy modeling [23, 20], often with Fokker–Planck-inspired objectives to align equilibrium and stationary laws [30]. Our three propagators Koopman, autoregressive NN, and score-guided Langevin, are representative of these families but are all evaluated inside a single, fixed latent space.
Coordinate-space generative simulators and MD surrogates.
Many recent models operate directly in Cartesian or SE(3)-equivariant coordinate space: GeoTDM uses diffusion for trajectory synthesis [12], F3low applies flow matching to coarse-grained rollouts [23], EGNO targets multi-step 3D dynamics with equivariant neural operators [50], and NeuralMD couples group-symmetric differential equations with multi-resolution protein–ligand dynamics [25]. Geometric diffusion models such as GeoDiff and DiffDock focus on static structures and poses [49, 7], while methods like LAST, ConfRover, and NeuralMD use learned representations to steer or unify MD-like dynamics [38, 34, 25]. Our setting is closer to system-specific surrogates such as LSS, GeoTDM, and ITO [35, 12, 32]: GLDP is trained per system to accelerate sampling of its equilibrium and kinetics, but keeps the all-atom decoder frozen and swaps only the propagator, enabling a controlled comparison of dynamical assumptions within a fixed latent representation.
Benchmark systems for protein dynamics.
[…]
@@ 3.1 GLDP: Propagation in the LD-FPG Latent Space
We adopt the LD-FPG encoder–decoder backbone [33]. The workflow follows a strict four-step sequence:
  1. 1.
    Encoding: A Chebyshev Graph Neural Network (ChebNet) maps all-atom coordinates $X(t)\in\mathbb{R}^{N\times 3}$ to per-atom embeddings.
  2. 2.
    Pooling: A deterministic pooling layer aggregates these embeddings into a single compact latent vector $z(t)\in\mathbb{R}^{d}$.
  3. 3.
    Propagation: The GLDP advances the latent state $z_{t}\to z_{t+1}$ using the learned transition function.
  4. 4.
    Decoding: The fixed decoder maps the evolved latent $z_{t+1}$ back to all-atom coordinates $\hat{X}(t+1)$.
The encoder produces a time series $\{z_{0},z_{1},\dots,z_{M}\}$ that captures conformational variation. We then learn a latent-time update of the form
@@ 3.4 Score-Guided Langevin Propagator
Our propagators evolve only the latent coordinate $z$ (no momenta, no barostat/box dynamics). The objective is to reproduce the configuration-space equilibrium of the reference MD at its thermodynamic state. Let $p_{X}(x)$ denote the equilibrium distribution of all-atom coordinates $x$ from the source MD, and let $E[\cdot]$ and $D[\cdot]$ be the fixed LD-FPG encoder and decoder. The target latent law is the pushforward $p_{Z}\;=\;E_{\#}\,p_{X}\quad\text{with}\quad z=E(x)$ and decoded samples are $\hat{x}=D(z)$. We fit propagators so that the stationary decoded trajectory $\{\hat{x}(t)\}$ matches reference statistics in $X$-space (e.g., dihedral histograms, contacts, and TM-distance free-energy surfaces). Because no Hamiltonian with momenta is integrated in latent space, we do not claim NVE energy conservation or exact physical kinetics. The score-guided Langevin propagator targets a canonical law in latent space whose invariant density is $p_{Z}$ when the score is exact; Koopman and the autoregressive NN model approximate the one-step transition kernel $K(z_{t+1}\!\mid z_{t})$ observed from $(z_{t},z_{t+1})$ under the fixed encoder. Unless stated otherwise, comparisons of timescales are diagnostic rather than calibrated to wall-clock time.
Rationale for Latent Dynamics.
Simulating directly in high-dimensional Cartesian space ($3N$) is computationally demanding and highly sensitive to error accumulation, which often violates steric constraints. GLDP circumvents this by propagating dynamics on a compressed manifold ($d\ll 3N$). Because the LD-FPG latent space encodes deformations relative to a physical reference structure, the dynamics are inherently centered within a valid conformational basin. This effectively decouples the simulation challenge: the fixed Decoder enforces local geometric validity (bond lengths and angles), allowing the Propagator to focus exclusively on learning slow, collective motions on a smoothed energy landscape free of high-frequency noise.
@@ 4 Results and Discussion
Key result. GLDP operates reliably across scales (ADP $\rightarrow$ 7JFL $\rightarrow$ 1R6W $\rightarrow$ $_{\mbox{1}}A$R) and reproduces the activation surface of $_{\mbox{2}}A$R; this cross scale pattern is the main message of the paper. See Appendix A.8 for $_{\mbox{1}}A$R ablations (encoder width, pooling size, and Langevin/denoiser settings).
@@ 4.1 Comparative Benchmarks on Trajectory Generation
To benchmark GLDP, we compared it against a diverse set of state-of-the-art models: the conceptually similar Latent Space Simulator (LSS) [35], the video-synthesis-based MD-Gen [16], and GeoTDM [12], a leading coordinate-based diffusion model. All methods were evaluated on equilibrium fidelity, flexibility, and—crucially—temporal kinetics across systems of increasing complexity: the canonical alanine dipeptide (ADP), the $\alpha$-helical 7JFL, the mixed $\alpha/\beta$ HIV-1 Protease (1R6W), and the functionally complex $_{\mbox{1}}A$R GPCR. The quantitative results are summarized in Table 1.111GeoTDM on 7JFL, 1R6W and $_{\mbox{1}}A$R exceeded 80 GB GPU memory of H100. MD-Gen does not support capped termini in our ADP setup; ADP results omitted.
Table 1: Benchmark of dynamics models. GLDP-NN is compared against LSS, MD-Gen, and GeoTDM. We include kinetic metrics (TICA timescales and decorrelation times) to assess temporal fidelity alongside structural ensemble metrics. Lower JSD is better. Higher Contact Correlation is better. For Kinetics, values closer to the Ground Truth are better. Bold indicates the best result among models. N/A: not applicable.
JSD Kinetics (steps)
System Model Dihedral (BB/SC) Coord. (BB/SC) Contact Corr. TICA ($t_{1}$) Decorr. ($t_{\text{decorr}}$)
ADP Ground Truth N/A N/A N/A 16.41 49
LSS 0.016 / N/A 0.621 / 0.633 N/A 15.80 46
GeoTDM 0.123 / N/A 0.444 / 0.395 N/A 12.20 35
GLDP 0.061 / N/A 0.366 / 0.386 N/A 10.14 28
7JFL Ground Truth N/A N/A N/A 660.34 2224
LSS 0.027 / 0.097 0.691 / 0.682 0.935 610.50 2050
MD-Gen 0.035 / 0.090 0.377 / 0.361 0.960 480.20 1600
GLDP 0.037 / 0.081 0.426 / 0.459 0.949 226.40 850
1R6W Ground Truth N/A N/A N/A 1670.11 2207
LSS 0.035 / 0.135 0.620 / 0.610 0.885 102.50 210
MD-Gen 0.050 / 0.105 0.190 / 0.210 0.980 450.30 980
GLDP 0.045 / 0.092 0.220 / 0.185 0.972 980.10 1650
$_{\mbox{1}}A$R Ground Truth N/A N/A N/A 1058.85 1281
LSS 0.134 / 0.146 0.657 / 0.602 0.971 Unstable Unstable
MD-Gen 0.033 / 0.088 0.106 / 0.117 0.941 85.40 150
GLDP 0.019 / 0.067 0.063 / 0.087 0.985 650.20 520
@@
We assess equilibrium fidelity via Jensen–Shannon divergence (JSD) over global coordinates and local dihedrals. On the small ADP system, LSS leads on backbone dihedrals, while GLDP attains the best coordinate JSD. On 7JFL, results are split, with LSS and MD-Gen performing well on backbone metrics, while GLDP best captures side-chain packing. The introduction of 1R6W challenges the models with a mixed topology; here, LSS captures local backbone dihedrals best and MD-Gen excels in backbone coordinate placement, yet GLDP achieves the lowest side-chain divergence (JSD 0.092 vs 0.135 for LSS) and superior side-chain coordinate fidelity (0.185). On the large $_{\mbox{1}}A$R, GLDP dominates, achieving the lowest JSDs for both coordinates and dihedrals, indicating the most faithful recovery of tertiary structure in large, membrane-bound systems. While Table 1 quantifies distributional error via JSD, we provide a visual comparison of the backbone and side-chain free-energy landscapes for ADP and $_{\mbox{1}}A$R against baselines (LSS, MD-Gen, GeoTDM) in App. Figure 6– 8.
@@
As a coarse proxy for structural stability, we compare the Pearson correlation between model and reference time-averaged contact maps. On 7JFL, MD-Gen best matches the contacts ($r=0.960$), followed by GLDP ($r=0.949$). This trend continues for the mixed-topology 1R6W, where MD-Gen ($r=0.980$) and GLDP ($r=0.972$) both maintain excellent tertiary definition, whereas LSS degrades ($r=0.885$), failing to preserve the $\beta$-sheet interface. However, on the most complex system, $_{\mbox{1}}A$R, GLDP attains the highest correlation ($r=0.985$), suggesting that while MD-Gen is strong on globular proteins, GLDP offers superior stability for complex transmembrane bundles.
Temporal Kinetics (TICA):
To verify that models capture dynamics rather than just independent samples, we quantify the slowest TICA timescale ($t_{1}$) and decorrelation time ($t_{\text{decorr}}$) (for more details check Appendix  A.7). On simpler systems (ADP, 7JFL), baselines like LSS effectively recover slow modes. However, complexity reveals a sharp divergence. On 1R6W, GLDP recovers a timescale of 980 steps and deep memory ($t_{\text{decorr}}=1650$), significantly outperforming the faster decorrelation of MD-Gen ($t_{\text{decorr}}=980$). Most notably on $_{\mbox{1}}A$R, LSS exhibits numerical instability (negative eigenvalues), whereas GLDP remains physically robust ($t_{1}\approx 650$ steps), proving it models coherent kinetics where baselines diverge.
@@
GLDP offers the strongest balance of structural realism and dynamic fidelity. While baselines like MD-Gen or LSS can excel on specific static metrics for simpler topologies, GLDP consistently outperforms on side-chain packing and flexibility for complex systems like 1R6W and $_{\mbox{1}}A$R. Crucially, the kinetic analysis confirms that GLDP is the only method stable enough to recover physically interpretable timescales on large GPCRs, positioning it as a highly dependable latent-space simulator.